# Can Direct Latent Model Learning Solve Linear Quadratic Gaussian Control?

### v2

@@ 5 Concluding remarks
We examined cost-driven state representation learning methods in time-varying LQG control. With a finite-sample analysis, we showed that a direct, cost-driven state representation learning algorithm effectively solves LQG. In the analysis, we revealed the importance of using multi-step cumulative costs as the supervision signal, and a separation of the convergence rates before and after step $\ell$, the controllability index, due to early-stage insufficient excitement of the system. A major limitation of our method is the use of history-based state representation functions; recovering the recursive Kalman filter would be ideal.
This work has opened up many opportunities for future research. An immediate question is how our cost-driven state representation learning approach performs in the infinite-horizon LTI setting. Moreover, one may wonder about the extent to which cost-driven state representation learning generalizes to nonlinear observations or systems. Investigating the connection with reduced-order control is also an interesting question, which may reveal the unique advantage of cost-driven state representation learning. Finally, one argument for favoring latent-model-based over model-free methods is their ability to generalize across different tasks; cost-driven state representation learning may offer a perspective to formalize this intuition.
Acknowledgement
YT, SS acknowledge partial support from the NSF BIGDATA grant (number 1741341). KZ’s work was mainly done while at MIT, and acknowledges partial support from Simons-Berkeley Research Fellowship. The authors also thank Xiang Fu, Horia Mania, and Alexandre Megretski for helpful discussions.
Appendix A Proposition on multi-step cumulative costs
The following proposition does not appear in the main body, but is important for analyzing CoReL (Algorithm 2) in Section 4.1.
Proposition 3.
Let $(z^{\ast\prime}_{t})_{t=0}^{T}$ be the state estimates by the Kalman filter under the normalized parameterization. If we apply $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$ for all $10\leq t\leq T-$ 1, then for $0\leq t\leq T$,
$\overline{c}_{t}:=c_{t}+c_{t+1}+\ldots+c_{t+k-1}=\|z^{\ast\prime}_{t}\|^{2}+\sum\nolimits_{\tau=t}^{t+k-1}\|u_{\tau}\|_{R^{\ast}_{\tau}}^{2}+b^{\prime}_{t}+e^{\prime}_{t},$
where $k=1$ for $10\leq t\leq\ell-$ 1 and $k=m\wedge(T-t+1)$ for $\ell\leq t\leq T$, $b^{\prime}_{t}=\mathcal{O}(k)$, and $e^{\prime}_{t}$ is a zero-mean subexponential random variable with $\|e^{\prime}_{t}\|_{\psi_{1}}=\mathcal{O}(kd_{x}^{1/2})$.
Proof.
$\overline{c}_{t}=\;$ $\sum\nolimits_{\tau=t}^{t+k-1}c_{\tau}$
=\;= $\sum\nolimits_{\tau=t}^{t+k-1}\big{(}\|\Phi^{\prime}_{\tau,t}z^{\ast\prime}_{t}+j^{\prime}_{\tau,t}\|_{Q^{\ast\prime}_{\tau}}^{2}+\|u_{\tau}\|_{R^{\ast}_{\tau}}^{2}+b^{\prime}_{\tau}+e^{\prime}_{\tau}\big{)}$
=\;= $(z^{\ast\prime}_{t})^{\top}\Big{(}\sum\nolimits_{\tau=t}^{t+k-1}(\Phi^{\prime}_{\tau,t})^{\top}Q^{\ast\prime}_{t}\Phi^{\prime}_{\tau,t}\Big{)}z^{\ast\prime}_{t}+\sum\nolimits_{\tau=t}^{t+k-1}\|u_{\tau}\|_{R^{\ast}_{\tau}}^{2}$
=\;= $\|z^{\ast\prime}_{t}\|^{2}+\sum\nolimits_{\tau=t}^{t+k-1}\|u_{\tau}\|_{R^{\ast}_{\tau}}^{2}+\overline{b}_{t}+\overline{e}_{t},$
Appendix B Auxiliary results
@@
$\|\bigl{\langle}x,y\bigr{\rangle}_{Q}\|_{\psi_{1}}=\||\bigl{\langle}x,y\bigr{\rangle}_{Q}|\|_{\psi_{1}}\leq\|\|x\|\|Q\|_{2}\|y\|\|_{\psi_{1}}=\|Q\|_{2}\cdot\|\|x\|\|y\|\|_{\psi_{1}}.$
@@
$\|\|x\|\|_{\psi_{2}}=\|\|\Sigma_{x}^{1/2}g\|\|_{\psi_{2}}\leq\|\|\Sigma_{x}^{1/2}\|_{2}\|g\|\|_{\psi_{2}}=\|\Sigma_{x}^{1/2}\|_{2}\|\|g\|\|_{\psi_{2}}.$
The distribution of $\|g\|_{2}$ is known as $\chi$ distribution, and we know that $\|\|g\|\|_{\psi_{2}}=c^{\prime}d^{1/4}$ for an absolute constant $c^{\prime}>0$ (see, e.g., (Wainwright, 2019)). Hence, $\|\|x\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{x}\|_{2}^{1/2}$. Similarly, $\|\|y\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{y}\|_{2}^{1/2}$. Since $\|\|x\|\|y\|\|_{\psi_{1}}\leq\|\|x\|\|_{\psi_{2}}\|\|y\|\|_{\psi_{2}}$ (see, e.g., (Vershynin, 2018, Lemma 2.7.7)), we have
@@
Let $x,y$ be random vectors of dimensions $d_{x},d_{y}$, respectively, defined on the same probability space. Then, $\|\mathrm{Cov}([x;y])\|_{2}\leq\|\mathrm{Cov}(x)\|_{2}+\|\mathrm{Cov}(y)\|_{2}$.
@@
Let $\mathrm{Cov}([x;y])=DD^{\top}$ be a factorization of the positive semidefinite matrix $\mathrm{Cov}([x;y])$, where $D\in\mathbb{R}^{(d_{x}+d_{y})\times(d_{x}+d_{y})}$. Let $D_{x}$ and $D_{y}$ be the matrices consisting of the first $d_{x}$ rows and the last $d_{y}$ rows of $D$, respectively. Then,
$\mathrm{Cov}([x;y])=DD^{\top}=\;$ $[D_{x};D_{y}][D_{x}^{\top},D_{y}^{\top}]$
=\;= $\begin{bmatrix}D_{x}D_{x}^{\top}&D_{x}D_{y}^{\top}\\ D_{y}D_{x}^{\top}&D_{y}D_{y}^{\top}\end{bmatrix}.$
Hence, $\mathrm{Cov}(x)=D_{x}D_{x}^{\top}$ and $\mathrm{Cov}(y)=D_{y}D_{y}^{\top}$. The proof is completed by noticing that
$\|\mathrm{Cov}([x;y])\|_{2}=\|D^{\top}D\|_{2}=\;$ $\|[D_{x}^{\top},D_{y}^{\top}][D_{x};D_{y}]\|_{2}$
=\;= $\|D_{x}^{\top}D_{x}+D_{y}^{\top}D_{y}\|_{2}$
[…]
=\;= $\|\mathrm{Cov}(x)\|_{2}+\|\mathrm{Cov}(y)\|_{2}.$
Appendix C Certainty equivalent linear quadratic control
Lemma 9 (Rank deficient linear quadratic control).
Consider the finite-horizon time-varying linear dynamical system $x_{t+1}=A^{\ast}_{t}x_{t}+B^{\ast}_{t}u_{t}+w_{t}$ with unknown $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{T-1}$.
  • Assume that $(A^{\ast}_{t})_{t=0}^{T-1}$ is uniformly exponentially stable (Assumption 1).
  • Assume that $x_{0}\sim\mathcal{N}(0,\Sigma_{0})$, $w_{t}\sim\mathcal{N}(0,\Sigma_{w_{t}})$ for all $10\leq t\leq T-$ 1, are independent, where $\Sigma_{0}$ and $(\Sigma_{w_{t}})_{t=0}^{T-1}$ do not necessarily have full rank.
  • Instead of $x_{t}$, we observe $x^{\prime}_{t}=x_{t}+\delta_{x_{t}}$, where the Gaussian noise vector $\delta_{x_{t}}$ can be correlated with $x_{t}$ and satisfy $\|\mathrm{Cov}(\delta_{x_{t}})^{1/2}\|\leq\varepsilon$ for all $0\leq t\leq T$.
  • Let $\Sigma_{t}:=\mathrm{Cov}(x_{t})$ and $\Sigma^{\prime}_{t}:=\mathrm{Cov}(x^{\prime}_{t})$ under control $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$ for all $10\leq t\leq T-$ 1. Assume $\sigma_{\min}^{+}((\Sigma_{t})^{1/2})\geq\beta>0$ and $\sigma_{\min}^{+}((\Sigma^{\prime}_{t})^{1/2})\geq\theta>0$ for all $0\leq t\leq T$, and that $\theta=\Omega(\varepsilon)$ with at least $n^{-1/4}$ dependence on $n$.
  • Assume $(Q^{\ast}_{t})_{t=0}^{T},(\hat{Q}_{t})_{t=0}^{T}$ are positive definite with $\mathcal{O}(1)$ operator norms, and $\|\hat{Q}_{t}-Q^{\ast}_{t}\|_{2}\leq\varepsilon$ for all $0\leq t\leq T$.
Collect $n$ trajectories using $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$ for some $\sigma_{u}>0$ and $10\leq t\leq T-$ 1. Identify $(\hat{A}_{t},\hat{B}_{t})_{t=0}^{T-1}$ by SysId (Algorithm 3), which uses ordinary least squares and takes a minimum Frobenius norm solution. Let $(\hat{K}_{t})_{t=0}^{T-1}$ be the optimal controller in system $((\hat{A}_{t},\hat{B}_{t},R^{\ast}_{t})_{t=0}^{T-1},(\hat{Q}_{t})_{t=0}^{T})$. Then, there exists an absolute constant $a>0$, such that if $n\geq a(d_{x}+d_{u}+\log(1/p))$, with probability at least $1-p$, $(\hat{K}_{t})_{t=0}^{T-1}$ is $\epsilon$-optimal for solving the LQR problem $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$ for ϵ=italic-$\epsilon=$
$\mathcal{O}(a^{T}((1+\beta^{-1})d_{x}\varepsilon+d_{x}(d_{x}+d_{u}+\log(1/p))^{1/2}n^{-1/2})),$
where dimension-free constant $a>0$ depends on the system parameters.
@@ Appendix C Certainty equivalent linear quadratic control
First of all, we establish that $(K^{\ast}_{t})_{t=0}^{T-1},(\hat{K}_{t})_{t=0}^{T-1}$ are bounded. Since $(Q^{\ast}_{t})_{t=0}^{T},(\hat{Q}_{t})_{t=0}^{T}$ are positive definite, the system is fully cost observable. By (Zhang and Zhang, 2021), the optimal feedback gains $(K^{\ast}_{t})_{t=0}^{T-1}$ have operator norms bounded by dimension-free constants depending on the system parameters. Note that (Zhang and Zhang, 2021) studies the boundedness of the Kalman filter, which is dual to our optimal control problem, but their results carry over. One can also see the boundedness from the compact formulation (Zhang et al., 2021), as described in the proof of Lemma 10, using known bounds for RDE solutions (Komaroff, 1996). The same argument applies to $(\hat{K}_{t})_{t=0}^{T-1}$ if we can show $(\hat{A}_{t},\hat{B}_{t})_{t=0}^{T-1}$ have $\mathcal{O}(1)$ operator norms. SysId (Algorithm 3) identifies $[\hat{A}_{t},\hat{B}_{t}]$ by ordinary least squares for all $10\leq t\leq T-$ 1. By Claim 1, with a union bound for all $10\leq t\leq T-$ 1, as long as $n\geq 16(d_{x}+d_{u}+\log(T/p))$, with probability at least $1-4p$, $\|[\hat{A}_{t},\hat{B}_{t}]\|_{2}=\mathcal{O}(\|[A^{\ast}_{t},B^{\ast}_{t}]\|_{2})=\mathcal{O}(1)$. Hence, with the same probability, $(\hat{K}_{t})_{t=0}^{T-1}$ have operator norms bounded by dimension-free constants depending on the system parameters. In the following, we assume $\|K^{\ast}_{t}\|_{2},\|\hat{K}_{t}\|_{2}\leq\kappa$ for all $10\leq t\leq T-$ 1, where $\kappa$ is a dimension-free constant that depends on system parameters.
Since $\Sigma_{t}$ can be rank deficient, we cannot guarantee $\|\hat{A}_{t}-A^{\ast}_{t}\|_{2}$ is small. Instead, for all $10\leq t\leq T-$ 1, by Lemma 5,
$([\hat{A}_{t},\hat{B}_{t}]-[A^{\ast}_{t},B^{\ast}_{t}])\textnormal{diag}(\Sigma_{t}^{1/2},\sigma_{u}I)=\mathcal{O}(\delta),$
where $\delta:=(1+\beta^{-1})\varepsilon+(d_{x}+d_{u}+\log(1/p))^{1/2}n^{-1/2}$, and $\epsilon_{x},\epsilon_{y},d_{1},d_{2}$ in Lemma 5 correspond to $\varepsilon,\varepsilon,d_{x}+d_{u},d_{x}$ here, respectively. This implies that $\|(\hat{A}_{t}-A^{\ast}_{t})(\Sigma_{t})^{1/2}\|_{2}=\mathcal{O}(\delta)$ and $\|\hat{B}_{t}-B^{\ast}_{t}\|_{2}=\mathcal{O}(\delta)$ for all $10\leq t\leq T-$ 1.
Let $\Xi_{t}$ denote the covariance of $x_{t}$ under state feedback controller $(K_{t})_{t=0}^{T-1}$, where $\|K_{t}\|\leq\kappa$ for all $10\leq t\leq T-$ 1. We claim that there exists $b_{t}>0$, such that $\Xi_{t}\preccurlyeq b_{t}\Sigma_{t}$. We prove it by induction. At step 00, clearly, $\Xi_{0}=\Sigma_{0}$. Suppose $\Xi_{t}\preccurlyeq b_{t}\Sigma_{t}$. For step $t+1$,
$\Xi_{t+1}=\;$ $(A^{\ast}_{t}+B^{\ast}_{t}K_{t})\Xi_{t}(A^{\ast}_{t}+B^{\ast}_{t}K_{t})^{\top}+\Sigma_{w_{t}}$
[…]
Hence, for $b\geq 1$,
$b\Sigma_{t+1}-\Xi_{t+1}$ $\succcurlyeq(b-2b_{t})A^{\ast}_{t}\Sigma_{t}(A^{\ast}_{t})^{\top}+b\sigma_{u}^{2}B^{\ast}_{t}(B^{\ast}_{t})^{\top}-2B^{\ast}_{t}K_{t}\Xi_{t}K_{t}^{\top}(B^{\ast}_{t})^{\top}.$
To ensure $b_{t+1}\Sigma_{t+1}\succcurlyeq\Xi_{t+1}$, it suffices to take $b_{t+1}=\max\{2,2\sigma_{u}^{-2}\kappa^{2}\|\Sigma_{t}\|_{2}\}b_{t}$. Hence, $b_{t}\leq a_{0}^{t}$, where constant $a_{0}>0$ is dimension-free and depends on system parameters. Note that if $(K_{t})_{t=0}^{T-1}$ stabilizes $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{T-1}$, then $\|\Xi_{t}\|_{2}=\mathcal{O}(1)$, and the bound on $b_{t}$ can be improved to $\sigma_{\min}^{+}(\Sigma_{t})^{-1}\|\Xi_{t}\|_{2}$; so the exponential dependence on $t$ results from not knowing whether $(K_{t})_{t=0}^{T-1}$ stabilizes the system. By the definition of the operator norm, we have $\|(\hat{A}_{t}-A^{\ast}_{t})\Xi_{t}^{1/2}\|_{2}\leq\|(\hat{A}_{t}-A^{\ast}_{t})(b_{t}\Sigma_{t})^{1/2}\|_{2}=\mathcal{O}(b_{t}^{1/2}\delta)$ for all $10\leq t\leq T-$ 1.
Let $\hat{\Xi}_{t}$ denote the covariance of $x_{t}$ under state feedback controller $(K_{t})_{t=0}^{T-1}$ in the system $(\hat{A}_{t},\hat{B}_{t})_{t=0}^{T-1}$. Then, by definition,
$\;\bigl{\|}\Xi_{t+1}-\hat{\Xi}_{t+1}\bigr{\|}_{2}$
$=$ $\;\Bigl{\|}(A^{\ast}_{t}+B^{\ast}_{t}K_{t})\Xi_{t}(A^{\ast}_{t}+B^{\ast}_{t}K_{t})^{\top}+\Sigma_{w_{t}}-\Bigl{(}\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}\hat{\Xi}_{t}\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}^{\top}+\Sigma_{w_{t}}\Bigr{)}\Bigr{\|}_{2}$
$=$ $\;\Bigl{\|}(A^{\ast}_{t}+B^{\ast}_{t}K_{t})\Xi_{t}(A^{\ast}_{t}+B^{\ast}_{t}K_{t})^{\top}-\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}\Xi_{t}\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}^{\top}$
$\quad+\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}(\Xi_{t}-\hat{\Xi}_{t})\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}^{\top}\Bigr{\|}_{2}$
$=$ $\;\mathcal{O}\Bigl{(}\kappa b_{t}^{1/2}\Bigl{\|}(A^{\ast}_{t}+B^{\ast}_{t}K_{t})\Xi_{t}^{1/2}-\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}\Xi_{t}^{1/2}\Bigr{\|}_{2}\Bigr{)}+\mathcal{O}(1+\kappa^{2})\|\hat{\Xi}_{t}-\Xi_{t}\|_{2},$
where the operator norms of $A^{\ast}_{t},B^{\ast}_{t}$ are $\mathcal{O}(1)$ and the operator norms of $\hat{A}_{t},\hat{B}_{t}$ are $\mathcal{O}(1)$ with probability at least $1-4p$. Since
$\Bigl{\|}(A^{\ast}_{t}+B^{\ast}_{t}K_{t})\Xi_{t}^{1/2}-\bigl{(}\hat{A}_{t}+\hat{B}_{t}K_{t}\bigr{)}\Xi_{t}^{1/2}\Bigr{\|}_{2}=\;$ $\Bigl{\|}(A^{\ast}_{t}-\hat{A}_{t})\Xi_{t}^{1/2}+(B^{\ast}_{t}-\hat{B}_{t})K_{t}\Xi_{t}^{1/2}\Bigr{\|}_{2}$
=\;= $\mathcal{O}(\delta\kappa b_{t}^{1/2}),$
we have $\bigl{\|}\hat{\Xi}_{t+1}-\Xi_{t+1}\bigr{\|}_{2}=\mathcal{O}(\delta\kappa^{2}b_{t}+(1+\kappa^{2})\|\hat{\Xi}_{t}-\Xi_{t}\|_{2})$. Combining with $\hat{\Xi}_{0}=\Xi_{0}=\Sigma_{0}$ gives
$\|\hat{\Xi}_{t}-\Xi_{t}\|_{2}=\mathcal{O}(b_{t}(a_{1}+a_{1}\kappa^{2})^{t}\delta)=\mathcal{O}((a_{2}+a_{2}\kappa^{2})^{t}\delta)$
for some dimension-free constants $a_{1},a_{2}>0$ that depend on the system parameters.
Let $J((K_{t})_{t=0}^{T-1})$ denote the expected cumulative cost under state feedback controller $(K_{t})_{t=0}^{T-1}$ in system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$ and $\hat{J}((K_{t})_{t=0}^{T-1})$ the corresponding expected cumulative cost in system $((\hat{A}_{t},\hat{B}_{t},R^{\ast}_{t})_{t=0}^{T-1},(\hat{Q}_{t})_{t=0}^{T})$. Notice that
$J((K_{t})_{t=0}^{T-1})=\mathbb{E}\Bigl{[}\sum\nolimits_{t=0}^{T}c_{t}\Bigr{]}=$ $\;\mathbb{E}\Bigl{[}\sum\nolimits_{t=0}^{T}x_{t}^{\top}(Q^{\ast}_{t})^{\prime}x_{t}\Bigr{]}$
$=$ $\;\mathbb{E}\Bigl{[}\sum\nolimits_{t=0}^{T}\left\langle(Q^{\ast}_{t})^{\prime},x_{t}x_{t}^{\top}\right\rangle_{F}\Bigr{]}$
$=$ $\;\sum\nolimits_{t=0}^{T}\left\langle(Q^{\ast}_{t})^{\prime},\Xi_{t}\right\rangle_{F},$
where $(Q^{\ast}_{t})^{\prime}:=(Q^{\ast}_{t}+K_{t}^{\top}R^{\ast}_{t}K_{t})$ for $10\leq t\leq T-$ 1 and $(Q^{\ast}_{T})^{\prime}:=Q^{\ast}_{T}$. Similarly, $\hat{J}((K_{t})_{t=0}^{T-1})=\sum_{t=0}^{T}\left\langle\hat{Q}^{\prime}_{t},\hat{\Xi}_{t}\right\rangle_{F}$, where $\hat{Q}^{\prime}_{t}=\hat{Q}_{t}+K_{t}^{\top}R^{\ast}_{t}K_{t}$ for $10\leq t\leq T-$ 1 and $\hat{Q}^{\prime}_{T}=\hat{Q}_{T}$. Hence,
$\big{|}J((K_{t})_{t=0}^{T-1})-\hat{J}((K_{t})_{t=0}^{T-1})\big{|}=\;$ $\sum\nolimits_{t=0}^{T}\left\langle(Q^{\ast}_{t})^{\prime}-\hat{Q}^{\prime}_{t},\Xi_{t}\right\rangle_{F}+\sum\nolimits_{t=0}^{T}\left\langle\hat{Q}^{\prime}_{t},\Xi_{t}-\hat{\Xi}_{t}\right\rangle_{F}$
=\;= $\varepsilon d_{x}\sum\nolimits_{t=0}^{T}\mathcal{O}(b_{t})+\kappa^{2}d_{x}\sum\nolimits_{t=0}^{T}\mathcal{O}((a_{2}+a_{2}\kappa^{2})^{t}\delta)$
=\;= $\varepsilon d_{x}\sum\nolimits_{t=0}^{T}\mathcal{O}(a_{0}^{t})+\delta\kappa^{2}d_{x}\sum\nolimits_{t=0}^{T}\mathcal{O}((a_{2}+a_{2}\kappa^{2})^{t})$
=\;= $\mathcal{O}(\delta d_{x}a^{T}),$
where $\|\hat{Q}^{\prime}_{t}\|_{2}=\mathcal{O}(\kappa^{2})$, $\varepsilon$ is absorbed into $\delta=(1+\beta^{-1})\varepsilon+(d_{x}+d_{u}+\log(1/p))^{1/2}n^{-1/2}$, and dimension-free constant $a>0$ depends on the system parameters.
Finally, let $(K^{\ast}_{t})_{t=0}^{T-1}$ be the optimal feedback gains in system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$. By the union bound, with probability at least $1-8p$,
$|J((\hat{K}_{t})_{t=0}^{T-1})-\hat{J}((\hat{K}_{t})_{t=0}^{T-1})|=\mathcal{O}(\delta d_{x}a^{T}),\quad|J((K^{\ast}_{t})_{t=0}^{T-1})-\hat{J}((K^{\ast}_{t})_{t=0}^{T-1})|=\mathcal{O}(\delta d_{x}a^{T}).$
Therefore,
$J((\hat{K}_{t})_{t=0}^{T-1})-J((K^{\ast}_{t})_{t=0}^{T-1})$
=\;= $J((\hat{K}_{t})_{t=0}^{T-1})-\hat{J}((\hat{K}_{t})_{t=0}^{T-1})+\hat{J}((\hat{K}_{t})_{t=0}^{T-1})-\hat{J}((K^{\ast}_{t})_{t=0}^{T-1})+\hat{J}((K^{\ast}_{t})_{t=0}^{T-1})-J((K^{\ast}_{t})_{t=0}^{T-1})$
=\;= $\mathcal{O}(d_{x}a^{T}\delta)$
=\;= $\mathcal{O}(a^{T}((1+\beta^{-1})d_{x}\varepsilon+d_{x}(d_{x}+d_{u}+\log(1/p))^{1/2}n^{-1/2})).$
The proof is completed by rescaling $8p$ to $p$. ∎
If the input of linear regression has full-rank covariance, then by Lemma 6, the system parameters can be fully identified. The certainty equivalent optimal controller in this case has a much better guarantee compared to the rank-deficient case, as shown in the following lemma.
Lemma 10 (Linear quadratic control).
Assume the finite-horizon linear time-varying system
$((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$ is stabilizable. Let $(K^{\ast}_{t})_{t=0}^{T-1}$ be the optimal controller and $(P^{\ast}_{t})_{t=0}^{T}$ be the solution to RDE (2.4) of system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$. Let $(\hat{K}_{t})_{t=0}^{T-1}$ be the optimal controller and $(\hat{P}_{t})_{t=0}^{T}$ be the solution to RDE (2.4) of system $((\hat{A}_{t},\hat{B}_{t},R^{\ast}_{t})_{t=0}^{T-1},(\hat{Q}_{t})_{t=0}^{T})$, where $\|\hat{A}_{t}-A^{\ast}_{t}\|_{2}\leq\epsilon$, $\|\hat{B}_{t}-B^{\ast}_{t}\|_{2}\leq\epsilon$ and $\|\hat{Q}_{t}-Q^{\ast}_{t}\|_{2}\leq\epsilon$. Then there exists dimension-free constant $\epsilon_{0}>0$ with $\epsilon_{0}^{-1}$ depending polynomially on system parameters, such that as long as $\epsilon\leq\epsilon_{0}$, $\|\hat{P}_{t}-P^{\ast}_{t}\|_{2}=\mathcal{O}(\epsilon)$, $\|\hat{K}_{t}-K^{\ast}_{t}\|_{2}=\mathcal{O}(\epsilon)$ for all $t\geq 0$, and $(\hat{K}_{t})_{t=0}^{T-1}$ is $\mathcal{O}((d_{x}\wedge d_{u})T\epsilon^{2})$-optimal in system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$.

### v3

@@ 5 Concluding remarks
We examined the cost-driven state representation learning methods in time-varying LQG control. With a finite-sample analysis, we showed that a direct, cost-driven state representation learning algorithm effectively solves LQG. In the analysis, we revealed the importance of using multi-step cumulative costs as the supervision signal, and the dependence on $\ell$, the controllability index, due to early-stage insufficient excitement of the system. For the same reason, our policy suboptimality gap has a significantly worsened dependence on $\ell$, as the latent model can only be partially identified, and the learned latent model may not be stable. Hence, an immediate question is how we can learn a stable latent model to improve the dependence on $\ell$. A major limitation of our method is the use of history-based state representation functions; recovering the recursive Kalman filter would be ideal.
In Part II of this work, we will explore how the cost-driven state representation learning approach performs in the infinite-horizon LTI setting, as well as discuss more opportunities that this work has opened up for future research.
Acknowledgment
YT, SS acknowledge partial support from NSF CCF-2112665 (TILOS AI Research Institute). KZ acknowledges the support from the Simons-Berkeley Research Fellowship, the U.S. Army Research Office grant W911NF-24-1-0085, and the NSF CAREER Award 2443704. SS and KZ mainly worked on this project while at MIT. The authors also thank Xiang Fu, Horia Mania, and Alexandre Megretski for helpful discussions.
Appendix
@@ Acknowledgment
$\|\bigl{\langle}x,y\bigr{\rangle}_{Q}\|_{\psi_{1}}=\;$ $\||\bigl{\langle}x,y\bigr{\rangle}_{Q}|\|_{\psi_{1}}$
$\leq\;$ $\|\|x\|\|Q\|_{2}\|y\|\|_{\psi_{1}}=\|Q\|_{2}\cdot\|\|x\|\|y\|\|_{\psi_{1}}.$
@@
$\|\|x\|\|_{\psi_{2}}=\;$ $\|\|\Sigma_{x}^{1/2}g\|\|_{\psi_{2}}$
$\leq\;$ $\|\|\Sigma_{x}^{1/2}\|_{2}\|g\|\|_{\psi_{2}}=\|\Sigma_{x}^{1/2}\|_{2}\|\|g\|\|_{\psi_{2}}.$
The distribution of $\|g\|_{2}$ is known as $\chi$ distribution, and we know that $\|\|g\|\|_{\psi_{2}}=c^{\prime}d^{1/4}$ for an absolute constant $c^{\prime}>0$ (see, e.g., Wainwright (2019)). Hence, $\|\|x\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{x}\|_{2}^{1/2}$. Similarly, $\|\|y\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{y}\|_{2}^{1/2}$. Since $\|\|x\|\|y\|\|_{\psi_{1}}\leq\|\|x\|\|_{\psi_{2}}\|\|y\|\|_{\psi_{2}}$ (see, e.g., Vershynin (2018, Lemma 2.7.7)), we have
@@
Let $x,y$ be random vectors of dimensions $d_{x},d_{y}$, respectively, defined on the same probability space. Then, $\|\mathbb{E}[[x;y][x;y]^{\top}]\|_{2}\leq\|\mathbb{E}[xx^{\top}]\|_{2}+\|\mathbb{E}[yy^{\top}\|_{2}$.
@@
Let $\mathbb{E}[[x;y][x;y]^{\top}]=DD^{\top}$ be a factorization of the positive semidefinite matrix $\mathbb{E}[[x;y][x;y]^{\top}]$, where $D\in\mathbb{R}^{(d_{x}+d_{y})\times(d_{x}+d_{y})}$. Let $D_{x}$ and $D_{y}$ be the matrices consisting of the first $d_{x}$ rows and the last $d_{y}$ rows of $D$, respectively. Then,
$\mathbb{E}[[x;y][x;y]^{\top}]=DD^{\top}=[D_{x};D_{y}][D_{x}^{\top},D_{y}^{\top}]=\begin{bmatrix}D_{x}D_{x}^{\top}&D_{x}D_{y}^{\top}\\ D_{y}D_{x}^{\top}&D_{y}D_{y}^{\top}\end{bmatrix}$.
Hence, $\mathbb{E}[xx^{\top}]=D_{x}D_{x}^{\top}$ and $\mathbb{E}[yy^{\top}]=D_{y}D_{y}^{\top}$. The proof is completed by noticing that
$\|\mathbb{E}[[x;y][x;y]^{\top}]\|_{2}=\|D^{\top}D\|_{2}=\;$ $\|[D_{x}^{\top},D_{y}^{\top}][D_{x};D_{y}]\|_{2}$
$=\;$ $\|D_{x}^{\top}D_{x}+D_{y}^{\top}D_{y}\|_{2}$
$\leq\;$ $\|D_{x}^{\top}D_{x}\|_{2}+\|D_{y}^{\top}D_{y}\|_{2}=\|\mathbb{E}[xx^{\top}]\|_{2}+\|\mathbb{E}[yy^{\top}]\|_{2},$
completing the proof. ∎
Lemma 13.
Let $x$ and $y$ be random vectors defined on the same probability space. Then, $\|\mathbb{E}[xy^{\top}]\|_{2}^{2}\leq\|\mathbb{E}[xx^{\top}]\|_{2}\cdot\|\mathbb{E}[yy^{\top}]\|_{2}$.
@@
Let $d_{x},d_{y}$ be the dimensions of the values of $x,y$, respectively. For any vectors $v\in\mathbb{R}^{d_{x}}$, $w\in\mathbb{R}^{d_{y}}$, by the Cauchy-Schwarz inequality,
$(v^{\top}\mathbb{E}[xy^{\top}]w)^{2}=\;$ $(\mathbb{E}[v^{\top}xy^{\top}w])^{2}$
$\leq\;$ $\mathbb{E}[(v^{\top}x)^{2}]\cdot\mathbb{E}[(w^{\top}y)^{2}]$
$=\;$ $\mathbb{E}[v^{\top}xx^{\top}v]\cdot\mathbb{E}[w^{\top}yy^{\top}w]$
$=\;$ $(v^{\top}\mathbb{E}[xx^{\top}]v)\cdot(w^{\top}\mathbb{E}[yy^{\top}]w).$
Taking the maximum over $v$, $w$ on both sides subject to $\|v\|,\|w\|\leq 1$ gives
$\|\mathbb{E}[xy^{\top}]\|_{2}^{2}\leq\|\mathbb{E}[xx^{\top}]\|_{2}\cdot\|\mathbb{E}[yy^{\top}]\|_{2},$
which completes the proof. ∎
Lemma 14.
Let $x^{\ast}$ and $\delta$ be $d$-dimensional random vectors defined on the same probability space. Define $x:=x^{\ast}+\delta$. Then,
$\|\mathbb{E}[xx^{\top}]-\mathbb{E}[x^{\ast}(x^{\ast})^{\top}]\|_{2}\leq\|\mathbb{E}[\delta\delta^{\top}]\|_{2}+2\|\mathbb{E}[x^{\ast}(x^{\ast})^{\top}]\|^{1/2}_{2}\|\mathbb{E}[\delta\delta^{\top}]\|_{2}^{1/2}.$